OpenStudy (trumpdonald):

HELP OR MY MOM WILL KILL ME!!!! What is the equation in point-slope form of the line that passes through the points (7, 5) and (−4, −1) ? y−1=611(x−4) y+4=116(x+1) y+1=611(x+4) y+1=67(x+4)

jagr2713 (jagr2713):

Well lets find slope. Plug in our equation into the slope formula and tel me what you get y2-y1 / x2-x1

OpenStudy (decentnabeel):

The Equation is.. y=mx+b m is the Slope and b is the y-intercept so that first find the slope using the two point m=slope=(y2-y1)/(x2-x1)

jagr2713 (jagr2713):

When you do get the slope then we will have M = ? and pick either first or second so (7,5) Then the point slope formula y-y1 = m(x-x1)

OpenStudy (decentnabeel):

y2=-1 and y1=5 x2=-4 and x1=7 Plug the value @tTrumpDonald

OpenStudy (triciaal):

asker off in 6 minutes?

OpenStudy (decentnabeel):

oka @triciaal

jagr2713 (jagr2713):

Well seem like you're gone? So lets solve. -1-5= -6 -4-7 = -11 m= 6/11 and (-4,-1) Now plug into point slope formula y+1 = 6/11(x+4)

OpenStudy (decentnabeel):

@jagr2713 stop this asker is offline

OpenStudy (triciaal):

@TrumpDonald we are not here to do the work for you. If you really need help you should be here to learn something and not devalue the learning process.

jagr2713 (jagr2713):

Usually when the asker leaves i just finish the work so when they do come back they can have an idea of what went on :D Their internet could of crash etc etc.

OpenStudy (mathmale):

Trump: Do you have the slope now and understand where it came from? The point-slope form is y=mx+b. Substitute your value for the slope, m, into this equation. Then, choose ONE of the 2 given points. Subst. the x- and y-values from this point into the above equation. Solve for the y-intercept, b. Show your work.